OpenStudy (anonymous):

The surface area of a cylinder is 289 cm² and the height is 21 cm. Find the diameter.

OpenStudy (shane_b):

\[SA_{cylinder}=2 \pi (\frac{d}{2})^2 + 2 \pi (\frac{d}{2}) h\]Given what you know:\[289cm^2=2 \pi (\frac{d}{2})^2 + 2 \pi (\frac{d}{2}) 21cm\]Now just solve for d.

OpenStudy (anonymous):

still dont get it

OpenStudy (shane_b):

I gave you the equation with the values filled in. At this point, you just need to solve for d which requires a bit of algebra including a quadratic equation to solve at the end. At any rate, you should end up with a resulting diameter of ~4cm.